OpenStudy (anonymous):

Find two positive real numbers whose product is a maximum and whose sum is 130.

OpenStudy (shamil98):

x+y = 130 A = xy where A is the product take the first equation and solve for y y = 130 -x Substitute that in the second equation A = x(130-x) A = 130x - x^2 Then take the derivative of the product in order to determine the maximum area: Let A' = 0 0 = 130 - 2x -130= -2x x = 65 Substitute that value in the first equation. 65 + y = 130 y = 65 Hence, two positive real numbers whose sum is 50 and whose product is a maximum are 65 and 65.

OpenStudy (shamil98):

Well, the product is actually 65(65) = 4225 my bad.

OpenStudy (shamil98):

whose sum is 130* ...

OpenStudy (anonymous):

thank you :D